OpenStudy (ksaimouli):

3sin^2theta+7sinsintheta +2=0

OpenStudy (lalaly):

sinsin?

OpenStudy (anonymous):

I answered this already?

OpenStudy (ksaimouli):

would u anser again

OpenStudy (anonymous):

replace sin(x) with a varaible, then solve the resulting quadratic

OpenStudy (anonymous):

*sin(theta)

OpenStudy (anonymous):

after replacing sin theta with a, we get: 3a^2+7a+2=0

OpenStudy (ksaimouli):

after that

OpenStudy (anonymous):

well you gotta factor it, you should get: (a+2)(3a+1)=0

OpenStudy (anonymous):

Then solve for a ineach term: a+2=0----->a=-2 3a+1=0---->a=-1/3

OpenStudy (anonymous):

then replace a with sin(theta)

OpenStudy (lalaly):

but\[-1 \le \sin \theta \le 1\] so neglect the a=-2 and solve for theta in sin(theta)= -1/3

OpenStudy (anonymous):

Nice job @lalaly

OpenStudy (lalaly):

lol thankyou :$